OpenStudy (anonymous):

Indicate the equation of the line that is the perpendicular bisector of the segment with endpoints (4, 1) and (2, -5).

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

first find equation of line y=3x-11 now the perpendicular bisector passes hought point (3, -2)

OpenStudy (alexandervonhumboldt2):

now find it's eqaution: y=-1/3x+b -2=-1/3*3+b b=-1 so the eqaution is y=-1/3x-1

OpenStudy (anonymous):

i cant use /

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

y=-1/3x-1

OpenStudy (anonymous):

thnx